OpenStudy (anonymous):

i need the 3 points to graph 5x-4y=-18

OpenStudy (anonymous):

-4y = - (18 + 5x) => 4y = 18 + 5x y = (18 + 5x)/4 from there you can choose any number to be your x plug it in to find y

OpenStudy (anonymous):

choose 3 different x , find their y corespondents and from there you should be able to graph it with no problem. your x (s) can be anything, like 0,1,2,3,4.6 , whatever you want.

OpenStudy (anonymous):

k thanks